# Qingdao 266237, China

### v1

@@ 6 MPSU from Public-Key Techniques
In this section, we describe how to construct a PK-MPSU achieving both linear computation and linear communication complexity. The construction of our PK-MPSU is formally presented in Figure 13.
@@
Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{rot}})$-hybrid model, assuming the security of MKR-PKE scheme, where $P_{1}=P_{\textnormal{ld}}$.
The security of the above protocol is based on security of the underlying building blocks, batch ssPMT and ROT, along with the rerandomizable property and indistinguishable multiple encryptions of MKR-PKE. For the complete proof, refer to Appendix 0.E.3.
  • Complexity Analysis. We provide a comprehensive complexity analysis for our PK-MPSU protocol in Appendix 0.F.3. The total costs are summarized below:
    • Offline phase. The offline computation complexity per party is $O(\gamma mn\log n)$. The offline communication complexity per party is $O(t\lambda m\log(\gamma n/t))$. The offline round complexity is $O(1)$.
    • Online phase. The online computation complexity per party is $O(mn)$ symmetric-key operations and $O(mn)$ public-key operations. The onlne communication complexity per party is $O((\gamma+\lambda)mn)$. The online round complexity is $O(\log\gamma+m)$.
  • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$. An OKVS scheme $(\mathsf{Encode_{H}},\mathsf{Decode_{H}})$. A MKR-PKE scheme $\mathcal{E}=(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$.
@@
      Hashing to bin. For $1\leq j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
@@
        For $2\leq j\leq m,1\leq b\leq B$: $P_{j}$ defines a vector $\vec{c}_{j}$ and sets $c_{j}^{b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{j}^{b}))$. $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.
        For $2\leq i<j$: $P_{i}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{i}$ acts as receiver with input $e_{i,j}^{b}$ and $P_{j}$ acts as sender. $P_{i}$ receives $r_{j,i,e_{i,j}^{b}}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ receives $r_{j,i,0}^{b},r_{j,i,1}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ computes $u_{j,i,e_{j,i}^{b}}^{b}=r_{j,i,e_{j,i}^{b}}^{b}\oplus c_{j}^{b}$, $u_{j,i,e_{j,i}^{b}\oplus 1}^{b}=r_{j,e_{j,i}^{b}\oplus 1}^{b}\oplus\mathsf{Enc}(pk,\perp)$, then sends $u_{j,i,0}^{b},u_{j,i,1}^{b}$ to $P_{i}$.
@@
      Messages decryptions and shufflings. $P_{1}$ samples $\pi_{1}:[(m-1)B]\to[(m-1)B]$ and computes $\vec{\mathsf{ct}}_{1}^{\prime\prime}=\pi_{1}(\vec{{\mathsf{ct}}_{1}^{\prime}})$. $P_{1}$ sends $\vec{\mathsf{ct}}_{1}^{\prime\prime}$ to $P_{2}$.
@@
Remark 2
When instantiating our PK-MPSU framework with EC MKR-PKE, the element space has to be set as EC points accordingly, which may limit its usage in a wide range of applications. We argue that the resulting PK-MPSU protocol is still useful in scenarios that the element space is exactly EC points. We demonstrate this by building the first multi-party private ID protocol (as described in Appendix 0.G), in which $\Pi_{\mathsf{PK\text{-}MPSU}}$ for EC point elements is used as a core building block.
7 Theoretical Comparison
In this section, we compare the construction of mq-ssPMT in [LG23] (mq-ssPMT for short) and our alternative construction built on batch ssPMT (batch ssPMT for short) theoretically. Then we compare our two protocols with the recent works [LG23] (which represents the MPSU protocol with the best concrete performance) and [GNT23] (which represents the MPSU protocol with the best asymptotic complexity), respectively. We emphasize that both of the two protocols are not in the standard semi-honest model.
7.1 Comparison Between mq-ssPMT and Batch ssPMT
Liu and Gao adapt the SKE-based mq-RPMT in [ZCL+23] to construct mq-ssPMT. The most expensive part of their construction is the secret-shared oblivious decryption-then-matching (ssVODM) [ZCL+23] protocol, which is to implement a decryption circuit and a comparison circuit by the GMW protocol [GMW87]. In total, the ssVODM protocol requires $(T+l-\log n-1)n$ AND gates, where $l$ is the bit length of set elements, and $T$ is the number of AND gates in the SKE decryption circuit and is set to be considerably large ($\approx 600$) according to their paper.
The costs of the batch ssPMT of size $B=1.27n$ [^8] We use stash-less Cuckoo hashing [PSTY19] with 3 hash functions, where $B=1.27n$. consist of the costs of batch OPPRF of size $1.27n$ and the costs of $1.27n$ instances of ssPEQT, where the batch OPPRF can be instantiated by extremely fast specialized protocol and the ssPEQT is implemented by the GMW protocol. Moreover, the state-of-the-art batch OPPRF construction [CGS22, RS21, RR22] can achieve linear computation and communication complexity with respect to $n$ In the ssPEQT, the parties engage in $1.27(\gamma-1)n$ AND gates, where $\gamma$ is the output length of OPPRF. Note that in the typical setting where $n\leq 2^{24},l\leq 128,\gamma\leq 64$, we have $(T+l-\log n-1)n\gg 1.27(\gamma-1)n$, which means that the number of AND gates desired involved in batch ssPMT is far smaller than mq-ssPMT. Therefore, the construction built on batch ssPMT greatly reduces the dependency on general 2PC and significantly decreases both computation and communication complexity by a considerable factor, compared to the mq-ssPMT (For a finer-grained asymptotic analysis of mq-ssPMT and batch ssPMT, refer to Appendix 0.F.4 and Appendix 0.F.1).
7.2 Comparison Between LG Protocol and Our SK-MPSU
We provide a comprehensive complexity analysis for the state-of-the-art LG protocol in Appendix 0.F.4. In Table 2, we present a comparison of the theoretical computation and communication complexity for each party in both offline and online phases between LG protocol and our SK-MPSU.
Computation Communication
Offline Online Offline Online
Leader Client Leader Client
Table 2: Asymptotic communication (bits) and computation costs of LG protocol and our SK-MPSU in the offline and online phases. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$. $T$ is the number of AND gate in an SKE decryption circuit, $T\approx 600$. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. $t$ is the noise weight in dual LPN, $t\approx 128$.
As depicted in Table 2, our offline communication complexity is comparable to theirs, our offline computation complexity and online computation and communication complexity of the leader are superior to theirs. Nevertheless, our online computation and communication complexity of client seem to be surpassed by theirs, as ours are quadratic in the number of parties $m$, while their is linear. We argue that it is because that in order to achieve security against arbitrary collusion, our protocol introduces extra overhead (from mss-ROT). Even so, for most applications when $m$ is moderate, and our online computation and communication costs still outperform theirs. This shall give the credit to our speedup by replacing their mq-ssPMT with batch ssPMT.
7.3 Comparison Between [GNT23] and Our PK-MPSU
Computation Communication
Offline Online Offline Online
Ours $\gamma mn\log n$ $mn$ $t\lambda m\log n$ $(\gamma+\lambda)mn$
Table 3: Asymptotic communication (bits) and computation costs of [GNT23] and our PK-MPSU protocol in the offline and online phases. In the offline phase, the computation is composed of symmetric-key operations; In the online phase, the computation is composed of public-key operations since we ignore symmetric-key operations. $n$ is the set size. $m$ is the number of participants. $\lambda$ and $\sigma$ are computational and statistical security parameter respectively and $\lambda=128$, $\sigma=40$. $l$ is the bit length of input elements. $\gamma$ is the output length of OPPRF. $t$ is the noise weight in dual LPN, $t\approx 128$.
@@ 8 Performance Evaluation
In this section, we provide implementation details and experimental results for our SK-MPSU and PK-MPSU protocols. Most previous works [KS05, Fri07, BA12, GNT23] lack open-source codes. The work of [VCE22] shows fairly poor performance in comparison with the state-of-the-art LG protocol. Therefore, we only compare our works with LG protocol whose complete implementation is available on https://github.com/lx-1234/MPSU.
In the implementation of LG protocol, they designate one party to generate share correlations in the offline phase and store them as local files, so that other parties can read these files and consume these share correlations in the online phase [^9] Refer to the function ShareCorrelation::generate() in ShareCorrelationGen.cpp of the repository https://github.com/lx-1234/MPSU.git. This implementation is not faithful to the protocol specifications, which is to let the parties pairwise execute share translation protocols [CGP20]. It has two pitfalls: (1) Their code does not support distributed execution. (2) It would lead to serious information leakage. Concretely, if it is the leader who generates share correlations, then it can learn the party to whom each element in the union belongs (which is the same security problem addressed by the execution of multi-party secret shuffle as we mentioned in the technical overview). If it is a client who generates share correlations, then it can learn the union. Moreover, their implementation gives incorrect results in several test cases when the set size $n$ is quite large (cf. Table 4). As a result, their code cannot be considered as a correct or secure implementation for MPSU.
To conduct a fair comparison, we replace their implementation of share correlation generation with our correct one.
@@ 8.1 Experimental Setup
We run LG and our two protocols on Ubuntu 22.04 with a single Intel i7-13700 2.10 GHz CPU (16 physical cores) and 64 GB RAM. We emulate the two network connections using Linux tc command. In the LAN setting, the bandwidth is set to be 10 Gbps with 0.1 ms RTT latency. In the WAN setting, the bandwidth is set to be 400 Mbps with 80 ms RTT latency. We record the running time as the maximal time from protocol begin to end, including messages transmission time. We compute the communication costs of the leader as the sum of the data it sent and received. For a fair comparison, we stick to the following setting for all protocols being evaluated:
@@
  • We test the balanced scenario by setting all $m$ input sets to be of equal size. In LG protocol and our SK-MPSU, each party holds $n$ 64-bit strings. In our PK-MPSU, we assume that each party holds $n$ elements encoded as elliptic curve points (We provide two versions of the implementation of our PK-MPSU, including encoding elements in uncompressed form and in compressed form respectively).
  • For each party, we use $m-1$ threads to interact with all other parties simultaneously and 4 threads to perform share correlation generation (in LG protocol and our SK-MPSU), Beaver Triple generation (in all three), parallel SKE encryption (in LG protocol), ciphertext rerandomization and partial decryption (in our PK-MPSU).
@@ 8.3 Experimental Results
We conduct an extensive experiment for the numbers of parties $\{3,4,5,7,9,10\}$ and a wide range of set sizes $\{2^{6},2^{8},2^{10},2^{12},2^{14},2^{16},2^{18},2^{20}\}$ in the LAN and WAN settings. We compare the performance of the protocols from four dimensions: online and total running time, and online and total communication costs. The results of online and total running time are depicted in Table 4. The results of online and total communication costs are depicted in Table 5.
@@
Online time (seconds) Total time (seconds)

### v4

@@ 6 MPSU from Public-Key Techniques
    In this section, we describe how to construct a PK-MPSU achieving linear computation and linear communication complexity. The construction is formally presented in Figure 12.
@@
    Protocol $\Pi_{\mathsf{PK\text{-}MPSU}}$ securely implements $\mathcal{F}_{\mathsf{mpsu}}$ against any semi-honest adversary corrupting $t<m$ parties in the $(\mathcal{F}_{\mathsf{bssPMT}},\mathcal{F}_{\mathsf{rot}})$-hybrid model, assuming the rerandomizable property and indistinguishable multiple encryptions (cf. Appendix 0.B) of MKR-PKE scheme.
    For the complete proof, refer to Appendix 0.D.3. A comprehensive complexity analysis and an exhaustive comparison with [GNT23] can be found in Appendix 0.E.
    Remark 2
    When instantiating our PK-MPSU framework with EC MKR-PKE, the element space is set as EC points accordingly, which might limit its usage in applications. We also identified the same issue in [GNT23]. However, we argue that this limitation is not inherent in our PK-MPSU (nor to the protocol in[GNT23]), as it can be addressed by finding an elliptic curve with efficient encoding and decoding algorithms or by exploring alternative instantiations of MKR-PKE.
    Despite the limitation, our EC-element-space PK-MPSU remains useful in many scenarios. For instance, we present the first multi-party private ID protocol in Appendix 0.F, where our EC-element-space PK-MPSU MPSU serves as a core subprotocol. Our multi-party private-ID takes advantages of our PK-MPSU and achieves linear computation and communication complexity and low communication costs as well.
    • Parameters. $m$ parties $P_{1},\cdots,P_{m}$. Size $n$ of input sets. The bit length $l$ of set elements. Cuckoo hashing parameters: hash functions $h_{1},h_{2},h_{3}$ and number of bins $B$. A MKR-PKE scheme $\mathcal{E}=(\mathsf{Gen},\mathsf{Enc},\mathsf{ParDec},\mathsf{Dec},\mathsf{ReRand})$.
@@
        Hashing to bins. $P_{1}$ does $\mathcal{T}_{1}^{1},\cdots,\mathcal{T}_{1}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{1})$. For $1<j\leq m$, $P_{j}$ does $\mathcal{C}_{j}^{1},\cdots,\mathcal{C}_{j}^{B}\leftarrow\mathsf{Cuckoo}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$ and $\mathcal{T}_{j}^{1},\cdots,\mathcal{T}_{j}^{B}\leftarrow\mathsf{Simple}_{h_{1},h_{2},h_{3}}^{B}(X_{j})$.
@@
          For $2\leq j\leq m,1\leq b\leq B$: $P_{j}$ defines $\vec{c}_{j}$ and sets $c_{j}^{b}=\mathsf{Enc}(pk,\mathsf{Elem}(\mathcal{C}_{j}^{b}))$. $\mathsf{Elem}(\mathcal{C}_{j}^{b})$ denotes the element in $\mathcal{C}_{j}^{b}$.
          For $2\leq i<j$: $P_{i}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{i}$ acts as receiver with input $e_{i,j}^{b}$ and $P_{j}$ acts as sender. $P_{i}$ receives $r_{i,j}^{b}=r_{j,i,e_{i,j}^{b}}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ receives $r_{j,i,0}^{b},r_{j,i,1}^{b}\in\{0,1\}^{\lambda}$. $P_{j}$ computes $u_{j,i,e_{j,i}^{b}}^{b}=r_{j,i,e_{j,i}^{b}}^{b}\oplus c_{j}^{b}$, $u_{j,i,e_{j,i}^{b}\oplus 1}^{b}=r_{j,i,e_{j,i}^{b}\oplus 1}^{b}\oplus\mathsf{Enc}(pk,\perp)$, then sends $u_{j,i,0}^{b},u_{j,i,1}^{b}$ to $P_{i}$.
@@
        Messages decryptions and shufflings.
@@
          $P_{1}$ samples $\pi_{1}:[(m-1)B]\to[(m-1)B]$ and computes $\vec{\mathsf{ct}}_{1}^{\prime\prime}=\pi_{1}(\vec{{\mathsf{ct}}_{1}^{\prime}})$. $P_{1}$ sends $\vec{\mathsf{ct}}_{1}^{\prime\prime}$ to $P_{2}$.
        2. (b)
@@ 7 Performance Evaluation
    In this section, we provide the implementation details and experimental results for our works. For most previous works [KS05, Fri07, BA12, GNT23] do not provide open-source code, and the protocol in [VCE22] shows fairly inferior performance in comparison with the state-of-the-art LG, here we only compare our protocols with LG, whose implementation is available on https://github.com/lx-1234/MPSU.
    During our experiments, we identify several issues with the LG implementation:
    1. 1.
      The code is neither a complete nor a secure implementation for MPSU, as it lacks the offline share correlation generation of the multi-party secret-shared shuffle protocol [EB22]. Instead of adhering to the protocol specifications, the implementation designates one party to generate and store some “fake” share correlations as local files during the offline phase, while the other parties read these files and consume the share correlations in the online phase [^11] Refer to the function ShareCorrelation::generate() in ShareCorrelationGen.cpp. This results in two serious consequences:
      • The distributed execution is not supported.
      • The security of multi-party secret-shared shuffle is compromised, leading to information leakage.
      To conduct a fair and complete comparison, we integrated our correct version of the share correlation generation implementation into their code.
    2. 2.
      The code is not a correct implementation, as it produces incorrect results when the set size $n$ increases beyond a certain threshold. We figure out that this issue arises from the use of a fixed parameter [^12] The variable batchsize in line 30 of circuit/TripleGen.cpp for generating Beaver Triples in the offline phase, which leads to an insufficient number of Beaver Triples for the online phase when $n$ is large. We modified their code to ensure correct execution and mark the cases where the original code fails with *.
@@ 7.1 Experimental Setup
    We conduct our experiments on an Alibaba Cloud virtual machine with Intel(R) Xeon(R) 2.70GHz CPU (32 physical cores) and 128 GB RAM. We emulate the two network connections using Linux tc command. In the LAN setting, the bandwidth is set to be 10 Gbps with 0.1 ms RTT latency. In the WAN setting, the bandwidth is set to be 400 Mbps with 80 ms RTT latency. We measure the running time as the maximal time from protocol begin to end, including messages transmission time, and the communication costs as the total data that leader sent and received. For a fair comparison, we stick to the following settings for all protocols:
@@
    • We test the balanced scenario by setting all $m$ input sets to be of equal size. In LG and our SK-MPSU, each party holds $n$ 64-bit strings. In our PK-MPSU, each party holds $n$ elements encoded as EC points in compressed form.
    • Each party uses $m-1$ threads to interact simultaneously with all other parties and 4 threads to perform share correlation generation (in LG and our SK-MPSU), Beaver Triple generation (in all three), parallel SKE encryption (in LG), ciphertext rerandomization and partial decryption (in our PK-MPSU).
@@ 7.3 Experimental Results
    We conduct extensive experiments across various numbers of parties $\{3,4,5,7,9,10\}$ and a wide range of set sizes $\{2^{6},2^{8},2^{10},2^{12},2^{14},2^{16},2^{18},2^{20}\}$ in the LAN and WAN settings. The performance of protocols is evaluated from four dimensions: online and total running time, and online and total communication costs. The results for running time / communication costs are depicted in Table 2. As we can see in the table, our protocols outperform LG in all the case studies, even with enhanced security.
  • Online computation improvement. Our SK-MPSU achieves a 3.910.0×3.9-10.0$\times$3.9 - 10.0 × speedup compared to LG in the LAN setting, and a $\times1$ 2.1 × speedup compared to LG in the WAN setting. For example, in the online phase to compute the union of $2^{20}$-size sets among 3 parties, our SK-MPSU runs in 4.4 seconds (LAN) and 33.1 seconds (WAN), which is 5.8×5.8$\times$5.8 × and 1.64×1.64$\times$1.64 × faster than LG that takes 25.2 seconds (LAN) and 54.1 seconds (WAN), respectively.
  • Total computation improvement. Our SK-MPSU achieves a 1.27.8×1.2-7.8$\times$1.2 - 7.8 × speedup compared to LG in the LAN setting. Our PK-MPSU achieves a 1.24.0×1.2-4.0$\times$1.2 - 4.0 × speedup compared to LG in the WAN setting. For example, to compute the union of $2^{20}$-size sets among 3 parties, our SK-MPSU runs in 96.2 seconds overall (LAN), which is 6.4×6.4$\times$6.4 × faster than LG that takes 610.8 seconds. To compute the union of $2^{18}$-size sets among 9 parties, our PK-MPSU runs in 2304 seconds overall (WAN), which is 2.8×2.8$\times$2.8 × faster than LG that takes 6381 seconds.
  • Online communication improvement. The online communication costs of our SK-MPSU is 1.24.9×1.2-4.9$\times$1.2 - 4.9 × smaller than LG. For example, in the online phase to compute the union of $2^{20}$-size sets among 3 parties, our SK-MPSU requires 455.6 MB communication, which is a 2.0×2.0$\times$2.0 × improvement of LG that requires 917.0 MB communication.
  • Total communication improvement. The total communication costs of our PK-MPSU is 3.036.5×3.0-36.5$\times$3.0 - 36.5 × smaller than LG. For example, to compute the union of $2^{18}$-size sets among 9 parties, our PK-MPSU requires 960.1 MB communication, a 34.8×34.8$\times$34.8 × improvement of LG that requires 33360 MB.
    In conclusion, our protocols exhibit distinct advantages and are suitable for different scenarios:
    • Our SK-MPSU: This protocol excels in online performance. Additionally, it also outperforms other protocols in nearly all cases in the LAN setting. Notably, the online phase takes only $4.4$ seconds for 3 parties with sets of $2^{20}$ items each, and $7$ seconds for 5 parties with sets of $2^{20}$ items each in the LAN setting. Therefore, our SK-MPSU is the optimal choice when online performance is the primary concern, or in high-bandwidth networks.
    • Our PK-MPSU: This protocol has the best total performance in the WAN setting, so it is preferred when the total cost is crucial (e.g. in the small-sets setting), at the same time in low-bandwidth networks. Note that our PK-MPSU is the only protocol with the capability of running among 10 parties with sets of $2^{18}$ items each in our experiments, which sufficiently indicates the superiority of its linear complexities.
@@
    Online Total = Offline + Online